OpenStudy (anonymous):

TRIGONOMETRY: If the central angle theta =10rad, and the area of the sector formed is 25square units, find the AREA of the circle. TRIGONOMETRY: If the central angle theta =10rad, and the area of the sector formed is 25square units, find the AREA of the circle. @Mathematics

OpenStudy (ash2326):

theta =10 theta * radius^2=area of sector 10*radius^2=25 radius^2=2.5 area of circle=pi*r^2 =3.14*2.5 7.85 thers is some mistake is central angle 10 dgrees or radian

OpenStudy (ash2326):

reply kyragayle is central angle theta is in radians or degrees

OpenStudy (anonymous):

radians

OpenStudy (phi):

Your problems can be answered using proportions. For this one, imagine the angle were pi (180º). Then you would know it corresponds to 1/2 the circle.

OpenStudy (phi):

or, for this problem, the area of the sector is a fraction of the total area. the fraction is the ratio of the angle to 2pi (using rads) so \[ \frac{10}{2\pi}A= 25 \]

OpenStudy (phi):

make sense?

OpenStudy (anonymous):

whats your final answer?

OpenStudy (phi):

What's yours?

OpenStudy (anonymous):

my answer is thousand so i know im wrong

OpenStudy (phi):

I'll answer, if you try this problem first: theta= pi rads  area of the sector is 25. What is the area of the circle

OpenStudy (anonymous):

the area of a circle is \[A= 1/2 r^2\theta\] ryt?

OpenStudy (anonymous):

the equation i mean

OpenStudy (phi):

almost. But use common sense on this one.

OpenStudy (anonymous):

im so confused

OpenStudy (phi):

If my problem is a pizza, and you ate the sector I labelled 25, how much of the pizza did you eat?

OpenStudy (anonymous):

25

OpenStudy (phi):

and how much is left over?

OpenStudy (anonymous):

25?

OpenStudy (phi):

and how big was the pizza?

OpenStudy (anonymous):

if im going to add it up, the part of the pizza i ate and the left over its 50 i think

OpenStudy (phi):

Yes. Now try this:  the sector is 25 (this is a bigger pizza) how big is the total pizza?